Primo car rental agency charges ​$38 per day plus ​$0.45 per mile. Ultimo car rental agency charges ​$46 per day plus ​$0.95  per mile. Find the daily mileage for which the Ultimo charge is twice the Primo charge.
The mileage is nothing.

Answer by Theo(13342)

primo charges 38 dollars per day plus .45 per mile.
ultimo charges 46 dollars per day plus .95 per mile.
you want the number of miles per day that makes ultimo charge twice what primo charges.

let x equal the number of miles.

the formula to solve is:

46 + .95x = 2 * (38 + .45*x)
simplify to get:
46 + .95x = 76 + .90x
subtract .90x from both sides of the equation and subtract 46 from both sides of the equation and simplify to get:
.05x = 30
solve for x to get:
x = 600

when the daily mileage is 600, then ultimo will cost twice as much as primo.

for ultimo, the total charge is 46 + .95 * 600 = 616 dollars.
for primo, the total charge is 38 + .45 * 600 =308 dollars.

616/308 = 2 which means the cost of ultimo for the day is twice the cost of primo for the day.